OpenStudy (anonymous):

Hal's sister is 5 years older than Hal. The product of their ages is 456. How old are Hal and his sister?

OpenStudy (campbell_st):

well let Hal's age by x then his sister's age is x + 5 so that means \[x \times (x + 5) = 456\] distribute and solve the quadratic for x

OpenStudy (anonymous):

i did that and i got 5 +- 43

OpenStudy (anonymous):

Is that correct ? How would I change that to the form i need?

OpenStudy (campbell_st):

well you have \[x^2 + 5x - 456 = 0\] so it factors to (x + 24)(x - 19) = 0

OpenStudy (campbell_st):

so you only need the positive solution, which is Hal's age add5 years for his sister's age

OpenStudy (anonymous):

well, i attempted at the quadratic equation. but i don't understand as to how i got 5+or-43 . would that be 43+ -5?